OpenStudy (anonymous):

Using the specific heat capacity value that you determined in Part II of the lab, what is the most probable identity of the unknown metal that you examined? Please explain your reasoning in complete sentences. I got .0054 as the heat capacity and I have no clue what to do from there.

OpenStudy (anonymous):

@Somy

OpenStudy (somy):

is it new experiment?

OpenStudy (anonymous):

I answered this question Using the formula qunknown metal = m × c × ΔT, calculate the specific heat of the metal. Use the data from your experiment for the unknown metal in your calculation. Show ALL your work. Q= 25 x 4.18 x (29.1 – 25.2) q= 25 x 4.18 x 3.9 q= 407 407= 25.605 x c x (100.3 – 29.1) 407= 25.605 x c x 2918 407= 74715 x c 407/74715= 74715/74715 x c .0054 = c

OpenStudy (somy):

good job!!!!

OpenStudy (somy):

okay so what are the values given as specific heat capacities?

OpenStudy (somy):

what metals and what values?

OpenStudy (anonymous):

hehe thank you I was able to answer it with your help on a previous question :)

OpenStudy (anonymous):

this is all that was given Known specific heat values — Iron: 0.444 J/g°C; Zinc: 0.390 J/g°C; Copper: 0.385 J/g°C, Aluminum: 0.900 J/g°C

OpenStudy (somy):

alright the value that u got is too small let me recheck the calculation

OpenStudy (anonymous):

I thought so sigh

OpenStudy (somy):

found your mistake :)

OpenStudy (somy):

407= 25.605 x c x (100.3 – 29.1) 407= 25.605 x c x 2918 407= 74715 x c 407/74715= 74715/74715 x c .0054 = c

OpenStudy (somy):

2918 what is this?

OpenStudy (somy):

(100.3 – 29.1)= 71.2

OpenStudy (anonymous):

I keep getting 2918.73??

OpenStudy (somy):

how? something is wrong its subtraction r u multiplying?

OpenStudy (somy):

round the value of q to 408 J because the value u get is 407.55

OpenStudy (somy):

Q= mc Delta T lets make C as subject C= Q/ mDelta T

OpenStudy (somy):

Delta T for metal will be 100.3-29.1= 71.2 m= 25.605 Q= 408 J

OpenStudy (somy):

now put the values into the formula

OpenStudy (anonymous):

i got .2232

OpenStudy (somy):

yeah 0.2233 basically

OpenStudy (somy):

no the closest value to the one we are getting is Copper: 0.385 J/g°C

OpenStudy (somy):

are you getting it?

OpenStudy (anonymous):

yeah!

OpenStudy (somy):

good :)

OpenStudy (somy):

so what are u getting as error?

OpenStudy (anonymous):

42

OpenStudy (somy):

good job :))

OpenStudy (anonymous):

thank you!

OpenStudy (somy):

u r welcome :))

OpenStudy (anonymous):

I have one last question and I just need help understanding what they're asking me

OpenStudy (somy):

okaay :) np post a new q :)